OpenStudy (anonymous):

Divide the following polynomial.Write your answer in order of descending powers of x. (x^3 + y^3)/ (x - y)

OpenStudy (anonymous):

@amistre64 Help me with this one Wolfram Doesn't take

OpenStudy (anonymous):

it takes but doesn't match where im putting answers

OpenStudy (amistre64):

\[-\frac{2}{y-x}x^3-x^2-xy-y^2\]or \[\frac{2}{x-y}x^3-x^2-xy-y^2\]

OpenStudy (anonymous):

man you fight with that machine easy lol

OpenStudy (anonymous):

\[y^2-xy-x^2-x^3\frac{ 2 }{ y-x }\]

OpenStudy (anonymous):

for it should be like that look on screen

OpenStudy (amistre64):

thats ascending powers of x that way

OpenStudy (anonymous):

so im right?

OpenStudy (anonymous):

i started it doesn't fit -x^3 :/

OpenStudy (anonymous):

something wrong

OpenStudy (amistre64):

long hand we get x^2+xy +y^2 -------------------- x-y | x^3 +y^3 x^3-x^2y ---------- x^2y x^2y-xy^2 ----------- xy^2+y^3 xy^2-y^3 ---------- 2y^3

OpenStudy (amistre64):

\[x^2+xy+y^2+\frac{2y}{x-y}\]

OpenStudy (amistre64):

2y^3 on that last part

OpenStudy (amistre64):

synthetic gives us x^3+ 0+ 0 +y^3 0 y y^2 y^3 y 1 y y^2 2y^3 x^2 x \[x^2+xy+y^2+\frac{2y^3}{x-y}\]

OpenStudy (anonymous):

x^2+xy+y^2+[ ]2y/x-y their is still one spot left

OpenStudy (anonymous):

need to be filled

OpenStudy (anonymous):

nvm your right its just me putting symbols wrong lol

OpenStudy (amistre64):

and thats just using the generic paint program on windows

OpenStudy (anonymous):

@amistre64 nice painting :)

OpenStudy (amistre64):

thnx :)